A two-point source interference pattern is generated in a swimming pool. A piece of styrofoam, located on the second nodal line, is 12.0 m from one source and 20.0 m from the other source. One wave crest takes 2.0 s to travel the 35.0 m width of the pool. Find the speed, wavelength, and frequency of the waves


1
Expert's answer
2021-04-19T07:22:12-0400

Answer

Speed of wave

V"=\frac{35}{2}=17.5m/s"





Wavelength

"\lambda=\frac{12\times2}{3}=8m"


Frequency

"f=\frac{\lambda}{V}=\frac{8}{17.5}=0.45Hz."